OpenStudy (anonymous):

select the product of (-8x8y)(x9y5)

OpenStudy (jdoe0001):

\(\large {\bf (-8x8y)(x^9y^5)\implies (-8x^18y^1)(x^9y^5)\\ \quad \\-8\cdot 8\cdot x^1\cdot x^9\cdot y^1\cdot y^5 \implies -8\cdot 8\cdot x^{1+9}y^{1+5} }\)